OpenStudy (anonymous):

If given f(x) = x + 4 and g(x) = –2x2 + 3x – 1, what is (f – g)(x)? A. (f – g)(x) = –2x2 + 4x – 3 B. (f – g)(x) = –2x2 – 2x + 5 C. (f – g)(x) = 2x2 + 2x + 5 D. (f – g)(x) = 2x2 – 2x + 5 Please Help

OpenStudy (ranga):

(f – g)(x) = f(x) - g(x)

OpenStudy (anonymous):

ok so that would be\[(x+4)-(2x ^{2}+3x-1)\]

OpenStudy (ranga):

-2x^2

OpenStudy (anonymous):

ok so \[(f-g)(x)=2x ^{2}-2x+5?\]

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

great Thanks @ranga !

OpenStudy (ranga):

You are welcome.